OpenStudy (anonymous):

The value of s varies jointly with m and p. If s=16 when m=8 and p=1/2, what is the value of s when m=5 and p=3?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

s ~ kmp so 16=k*8*(1/2) ---> k=4 so s=4*m*p. now plug in those values of m, p to figure out s.

OpenStudy (anonymous):

So my answer is 60 =)

OpenStudy (anonymous):

correct... :)